a tennis court is 36 feet wide and 78 feet long. A fence is to be built around the court 10 feet from the edges of the court. How much fencing will be needed?
Answer by fcabanski(1391)

If you draw it, you will see the fence is 10 feet longer than the court on each side (98 feet long) and 10 feet wider than the court on each side (56 feet wide.) The perimeter is the length of fence needed. Perimeter is 2L + 2W = 2*56 + 2*98 = 112 + 196 = 308feet of fencing.